# A Selective Homomorphic Encryption Approach for Faster Privacy-Preserving Federated Learning

### v2

@@ Abstract
Federated learning is a machine learning method that supports training models on decentralized devices or servers, where each holds its local data, removing the need for data exchange. This approach is especially useful in healthcare, as it enables training on sensitive data without needing to share them. The nature of federated learning necessitates robust security precautions due to data leakage concerns during communication. To address this issue, we propose a new approach that employs selective encryption, homomorphic encryption, differential privacy, and bit-wise scrambling to minimize data leakage while achieving good execution performance. Our technique FAS (fast and secure federated learning) is used to train deep learning models on medical imaging data. We implemented our technique using the Flower framework and compared with a state-of-the-art federated learning approach that also uses selective homomorphic encryption. Our experiments were run in a cluster of eleven physical machines to create a real-world federated learning scenario on different datasets. We observed that our approach is up to 90% faster than applying fully homomorphic encryption on the model weights. In addition, we can avoid the pretraining step that is required by our competitor and can save up to 46% in terms of total execution time. While our approach was faster, it obtained similar security results as the competitor.
Index Terms:
@@
I Introduction
Federated learning, a technique introduced by Google in 2016, offers a way to train machine learning models across decentralized data from mobile devices [1]. This approach has since gained widespread interest in both academic research and industry applications. By allowing model training on data that remains locally stored, federated learning helps protect privacy, making it an attractive solution for sensitive data environments. For example, in healthcare, it enables institutions to build shared machine learning models without the need to transfer patient data, thereby maintaining compliance with data protection laws. Overall, federated learning facilitates secure, privacy-conscious model training across decentralized networks.
The utilization of federated learning in sensitive domains such as healthcare necessitates robust security precautions due to its reliance on internet connections and the inherent risks of unauthorized data access and potential leaks. To address these privacy concerns, various security techniques have been integrated into federated learning frameworks. This research focuses on the implementation and evaluation of cryptology-based security precautions in federated learning.
This paper introduces a detailed experimental evaluation of different security techniques applied to federated learning, focusing on their impact on performance, computational overhead, and privacy protection. Specifically, we examine the effectiveness of homomorphic encryption, differential privacy, and a custom-developed fast and secure homomorphic encryption (FAS) technique. These methods were tested using different image datasets in a federated learning scenario simulated in a cloud environment.
@@ I Introduction
Implementation of Security Techniques: We implemented secure aggregation methods, including homomorphic encryption, differential privacy, and FAS, within a federated learning model.
Development of FAS: We propose a novel FAS technique combining selective encryption, differential noise addition, and bitwise scrambling to achieve high security with minimal computational overhead.
@@
Performance Evaluation: We provide a comparative analysis of encryption techniques, highlighting their overhead, scalability, and resilience against model inversion attacks, using metrics like MSSIM and VIFP.
@@
Practical Insights: We assess the trade-offs between privacy, performance, and scalability, providing actionable insights for deploying secure federated learning systems in real-world applications.
The results demonstrate that FAS achieves a balance between robust security and computational efficiency, outperforming traditional techniques in scenarios where real-time performance and scalability are critical. The findings also highlight the suitability of FAS for resource-constrained environments like healthcare, where both privacy and efficiency are paramount.
The rest of this paper is organized as follows: The Datasets section describes the different image datasets used in our study. In Background on Deep Learning Models and Security Techniques, we discuss the deep learning models evaluated in our experiments, including ResNet-50 [2], DenseNet121 [3], EfficientNetB0 [4], MobileNet V2 [5], and elaborate on the security techniques applied.
The Experimental Setup for Federated Learning details our use of CloudLab [10] as the testbed and the Flower framework [11] for federated learning experiments. The Evaluation section presents encryption overhead and time costs across datasets and models, with a detailed time comparison of security techniques.
In Comparison of Security Techniques, we evaluate differential privacy [15], homomorphic encryption [16], and FAS, and compare these with similar works. The Practical Implications section discusses efficiency considerations and the balance between security and performance, followed by a Visual Comparison of FAS Approaches. Finally, we conclude with a summary of our findings and future research directions.
II Background and Related Work
@@
Privacy-preserving techniques in federated learning have gained attention due to the need to protect sensitive data across distributed devices. Federated learning, introduced by McMahan et al. [23], enables decentralized model training but poses risks of data leakage through model updates.
Homomorphic encryption (HE) offers a robust solution, allowing computations on encrypted data while maintaining confidentiality. Gentry’s fully homomorphic encryption (FHE) [16] made arbitrary encrypted computations feasible, though its high computational overhead limits practical use [17, 28].
II-A Efficiency Enhancements in Homomorphic Encryption
Several advancements have sought to improve the efficiency of homomorphic encryption:
Fan and Vercauteren’s BGV Scheme [25]: This scheme optimizes the bootstrapping process, making HE more practical by efficiently handling both addition and multiplication on ciphertexts. This approach has been critical in making HE viable for more complex computations in federated learning environments [33]. Xu et al. [41] proposed HybridAlpha, an efficient privacy-preserving federated learning system that optimizes both privacy and performance. In comparison to our FAS approach, HybridAlpha leverages different privacy mechanisms such as differential privacy and secure aggregation to ensure efficient federated learning while maintaining data confidentiality.
Chillotti et al.’s TFHE Scheme [26]: TFHE introduced a faster bootstrapping process, making it highly efficient for secure Boolean computations, thus expanding the potential use of FHE in time-sensitive environments [34].
Ducas and Micciancio’s FHEW Scheme [27]: FHEW further improved the speed of homomorphic operations by optimizing the bootstrapping process, making FHE more feasible for larger-scale applications [27].
Cheon et al.’s CKKS Scheme [29]: This scheme supports approximate arithmetic operations on encrypted data, making it particularly efficient for machine learning and statistical analysis tasks where exact precision is not always required [36].
Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme [28]: This work built on the BGV scheme, introducing further optimizations that reduce the computational overhead of FHE, particularly in noise management, which is crucial for maintaining the accuracy of federated learning models under encryption [28].
II-B Differential Privacy (DP) in Federated Learning
Differential privacy (DP) mitigates privacy risks by injecting controlled noise into model updates. Initially proposed by Dwork et al. [30], its integration into deep learning was further developed by Abadi et al. [31]. Subsequent work has refined DP methods to balance privacy and model accuracy [35].
II-C Selective Encryption in Federated Learning
Selective encryption reduces computational burden by encrypting only a portion of model parameters while maintaining security. Wu et al. [32] demonstrated its effectiveness in reducing overhead without compromising privacy. Similar approaches focus on optimizing encryption efficiency in federated learning [37, 38]. Zhang et al. [45] proposed BatchCrypt, an efficient homomorphic encryption framework that batches model updates to reduce the computational and communication overhead in federated learning. While BatchCrypt improves efficiency compared to traditional HE-based methods, it still incurs non-trivial overhead. Our approach further optimizes encryption by selectively encrypting a fraction of updates, achieving similar security guarantees with significantly lower computational costs. This paper extends these concepts by integrating selective encryption with differential noise and bitwise scrambling, reducing homomorphic encryption overhead while enhancing privacy protection.
II-D Bitwise Scrambling for Federated Learning Security
Bitwise scrambling further obfuscates partially encrypted data by rearranging bits using a cryptographic key. This technique complicates adversarial analysis and enhances security without significant computational cost.
@@ II-D Bitwise Scrambling for Federated Learning Security
Yang et al. [24] applied bitwise scrambling to secure data transmission, effectively preventing direct reconstruction of encrypted models. Halevi et al. [39] demonstrated its use with HE to resist chosen-ciphertext attacks while maintaining efficiency.
Combining selective encryption, differential privacy, and bitwise scrambling offers an effective privacy-preserving solution for federated learning, balancing security and computational efficiency for resource-constrained applications.
III Motivation and Threat Model
III-A Motivation
The adoption of federated learning (FL) in sensitive domains such as healthcare and finance necessitates robust privacy-preserving mechanisms. While FL mitigates data exposure by keeping raw data localized, model updates remain vulnerable to attacks like model inversion and gradient leakage, which can reveal sensitive information such as patient records or financial transactions.
Existing privacy-preserving techniques, such as fully homomorphic encryption (FHE) and differential privacy (DP), have limitations. FHE ensures confidentiality but incurs significant computational and communication overhead, making it impractical for resource-constrained settings. DP, though computationally efficient, can degrade model accuracy by introducing noise, which is problematic in high-precision applications.
To address these challenges, this work proposes the FAS approach, which balances security and efficiency by selectively encrypting critical model updates while applying differential noise and bitwise scrambling to the remaining data. This reduces computational overhead while maintaining strong privacy guarantees, making it particularly suitable for federated learning in time-sensitive applications like medical imaging and real-time analytics.
III-B Threat Model
@@
Federated learning systems are susceptible to privacy threats due to decentralized training and the exchange of model updates. A semi-honest adversary may follow the protocol but attempt to infer sensitive information from intercepted or compromised data.
@@
Adversaries can access model updates between clients and the central server, potentially compromising either to extract insights. Common attack vectors include gradient inversion to reconstruct training data, membership inference to verify specific data participation, and statistical analysis to infer private attributes. While SSL encryption secures transmissions, it does not prevent adversaries from analyzing transmitted updates.
@@
IV Proposed Model
This research introduces a novel FAS (fast and secure federated learning) approach tailored for secure and efficient federated learning. The model aims to strike a balance between robust privacy preservation and computational efficiency, addressing the limitations of fully homomorphic encryption (high overhead) and differential privacy (accuracy degradation). Our approach integrates three key components: selective encryption, differential noise addition, and bitwise scrambling, providing a multi-layered defense against privacy attacks while maintaining high performance.
IV-A Selective Encryption
Rather than encrypting the entire model, FAS selectively encrypts a fixed percentage of critical model parameters using a homomorphic encryption scheme. This subset is chosen uniformly, ensuring that sensitive components are protected without the computational burden of encrypting the entire dataset. By encrypting only a fraction of the data, FAS significantly reduces the overhead associated with full encryption while maintaining strong security for the most sensitive parameters.
IV-B Differential Noise Addition
For the remaining unencrypted parameters, differential noise is added to obfuscate their true values. The noise is carefully calibrated using a Laplacian mechanism, ensuring that individual data points cannot be inferred while preserving the overall utility of the model. This approach leverages the principles of differential privacy to provide an additional layer of protection for unencrypted data.
IV-C Bitwise Scrambling
To further enhance privacy, unencrypted parameters are subjected to bitwise scrambling using a cryptographic key. This process disrupts patterns in the data, making it computationally infeasible for adversaries to reconstruct the original values without access to the scrambling key. By combining scrambling with noise addition, the model ensures that unencrypted data is protected against both statistical analysis and direct reconstruction attacks.
IV-D Federated Learning Integration
The FAS model integrates seamlessly into the federated learning workflow. Each client encrypts and obfuscates its local model updates using the proposed techniques before transmitting them to the central server. The server aggregates the updates without requiring decryption, preserving privacy throughout the computation. The aggregated model is then securely shared with clients for the next training iteration.
IV-E Security and Efficiency Trade-Off
The proposed model achieves a balance between security and efficiency:
  • Security: By encrypting critical parameters and obfuscating the rest, FAS provides strong resistance to model inversion and membership inference attacks, with security results comparable to full homomorphic encryption.
  • Efficiency: The selective encryption strategy reduces computational and communication overhead by up to 90%, making it suitable for resource-constrained federated learning environments.
This layered approach ensures that sensitive data remains protected while enabling efficient model training, making FAS an effective solution for privacy-preserving federated learning in applications like healthcare and finance.
IV-F Privacy-Preserving Techniques
IV-F1 Homomorphic Encryption
Homomorphic encryption (HE) is a cryptographic technique that allows computations to be performed directly on encrypted data without decrypting it. This ensures that sensitive information remains confidential throughout the computation process.

### v4

@@ Abstract
Federated learning (FL) has come forward as a critical approach for privacy-preserving machine learning in healthcare, allowing collaborative model training across decentralized medical datasets without exchanging clients’ data. However, current security implementations for these systems face a fundamental trade-off: rigorous cryptographic protections like fully homomorphic encryption (FHE) impose prohibitive computational overhead, while lightweight alternatives risk vulnerable data leakage through model updates. To address this issue, we present FAS (Fast and Secure Federated Learning), a novel approach that strategically combines selective homomorphic encryption, differential privacy, and bitwise scrambling to achieve robust security without compromising practical usability. Our approach eliminates the need for model pretraining phases while dynamically protecting high-risk model parameters through layered encryption and obfuscation. We implemented FAS using the Flower framework and evaluated it on a cluster of eleven physical machines. Our approach was up to 90% faster than applying FHE on the model weights. In addition, we eliminated the computational overhead that is required by competitors such as FedML-HE and MaskCrypt. Our approach was up to 1.5$\times$ faster than the competitors while achieving comparable security results.
Experimental evaluations on medical imaging datasets confirm that FAS maintains similar security results to conventional FHE against gradient inversion attacks while preserving diagnostic model accuracy. These results position FAS as a practical solution for latency-sensitive healthcare applications where both privacy preservation and computational efficiency are requirements.
Keywords:
@@
1 Introduction
Federated learning (FL), first introduced by Google in 2016, enables machine learning models to be trained across decentralized datasets stored on distributed devices such as mobile phones federated . This method has gained widespread interest across academia and industry because of its effectiveness in safeguarding data privacy. By keeping training data localized and only sharing model updates, FL avoids direct data exposure, making it particularly valuable in regulated sectors like healthcare. For instance, hospitals can collaboratively train diagnostic models without transferring sensitive patient records, ensuring compliance with privacy regulations.
Nonetheless, the decentralized structure of FL brings about distinct security challenges. While data remains on local devices, the transmission of model parameters over networks creates vulnerabilities, including risks of unauthorized access or data leakage during exchange. Existing research often assumes that data localization inherently guarantees privacy, overlooking these communication-phase threats. Current security strategies in FL, such as encrypting communication channels or perturbing shared gradients, are frequently implemented in isolation, without systematic analysis of their combined impact on computational efficiency, model accuracy, or multi-layered privacy preservation.
Recent advances, such as FedML-HE jin2024fedmlheefficienthomomorphicencryptionbasedprivacypreserving and MASKCRYPT 10506637 , attempt to address these challenges through selective encryption. FedML-HE employs gradient sensitivity analysis during pre-training to identify critical weights for HE, but its reliance on client-specific masks introduces aggregation inconsistencies and computational overhead. MASKCRYPT optimizes gradient-guided masks to reduce communication costs, yet its security requires multiple rounds to stabilize, leaving early training phases vulnerable. Both methods incur significant computational penalties—FedML-HE from pre-training and MASKCRYPT from per-round mask recalibration—while struggling to balance real-time efficiency with robust privacy guarantees. These constraints point to the necessity of a cohesive solution that removes the need for initialization phases, reduces computational effort per round, and maintains stable security throughout all FL rounds.
Motivated by the aforementioned reasons, we propose FAS (Fast and Secure Homomorphic Encryption), a novel method designed to address these gaps by unifying cryptology-based security mechanisms tailored for federated architectures. Unlike FedML-HE and MASKCRYPT, FAS eliminates pre-training phases and per-round mask recalibration, instead combining selective homomorphic encryption with noise-and-scrambling mechanisms to secure a fixed percentage of model weights. This ensures consistent privacy guarantees from the first training round while minimizing computational overhead. We present a comparative study of FAS against conventional HE, differential privacy, and state-of-the-art baselines (FedML-HE and MASKCRYPT). Through simulations across cloud-based federated environments and diverse medical imaging datasets (e.g., Kidney, Lung, COVID), we rigorously measure how each method balances privacy protection, computational efficiency, and model utility.
To validate FAS’s efficacy, we conducted extensive experiments across medical imaging tasks and model architectures. Our evaluation shows that FAS offers high efficiency comparable to leading methods, all while preserving strong privacy protections. On datasets like Kidney and COVID, FAS reduced encryption overhead by up to 90% compared to FHE, with training times as low as 52 minutes for MobileNetV2 (vs. 610 minutes for full encryption). Crucially, FAS eliminated the pre-training and mask recalibration costs inherent to FedML-HE and MASKCRYPT, achieving 1.5×× faster execution than these baselines (e.g., 69 vs. 99 minutes for EffNetB0 on Kidney data). Security evaluations using MSSIM and VIFP confirmed that even at 10% encryption, FAS effectively thwarts model inversion attacks, with MSSIM scores of 58%—outperforming FedML-HE (52%) and MASKCRYPT (55%) in early-round privacy. FAS’s VIFP scores stabilized within 5 rounds, contrasting with MASKCRYPT’s delayed convergence. This combination of speed and robustness positions FAS as a practical solution for latency-sensitive federated deployments, where traditional methods trade excessive computation for marginal security gains.
@@ 1 Introduction
  • Implementation of Security Techniques: We implement three security mechanisms—HE gentry2009 , differential privacy differential , and the proposed FAS technique—within a FL framework, demonstrating their operational feasibility.
  • Development of FAS: We design a novel Fast and Secure Homomorphic Encryption method that integrates selective encryption with noise injection and bitwise scrambling to enhance security while maintaining computational efficiency.
@@
  • Performance Evaluation: We conduct a comprehensive comparison of encryption techniques using standardized metrics (MSSIM, VIFP) to assess their computational overhead, scalability, and resistance against model inversion attacks.
@@
  • Practical Insights: We characterize the trade-offs between privacy guarantees and system performance, offering implementable guidelines for deploying secure FL in real-world scenarios.
    Experimental results demonstrate that FAS achieves superior balance between security and efficiency compared to conventional methods, particularly in latency-sensitive applications like healthcare systems where privacy and responsiveness are critical.
2 Background and Related Work
@@
Privacy-preserving techniques in FL have gained significant attention due to the critical need to protect sensitive data across distributed devices. FL, initially introduced by McMahan et al. mcmahan2017communication , enables decentralized model training by sharing parameter updates instead of raw data. However, this approach introduces inherent risks of data leakage through exposed model gradients, necessitating robust cryptographic safeguards to prevent adversarial reconstruction of private datasets.
2.1 Efficiency Enhancements in Homomorphic Encryption
HE has emerged as a foundational technology for secure computation on encrypted data. Gentry’s pioneering FHE scheme gentry2009 first demonstrated the feasibility of arbitrary computations on ciphertexts, though its substantial computational overhead limited practical adoption. Subsequent research has focused on optimizing efficiency:
Fan and Vercauteren’s BGV Scheme fan2012somewhat introduced optimizations to the bootstrapping process, enabling faster processing of encrypted data by supporting basic arithmetic functions such as summation and product computations. This advancement significantly improved the practicality of HE for complex computations in FL environments, as further validated by Smart and Vercauteren smart2014fully .
Chillotti et al.’s TFHE Scheme chillotti2016faster achieved faster bootstrapping for secure Boolean operations, making FHE viable for real-time applications requiring rapid iterative computations. Ducas and Micciancio’s FHEW Scheme ducas2015faster further accelerated homomorphic operations through bootstrapping refinements, broadening HE’s applicability to large-scale distributed systems.
Cheon et al.’s CKKS Scheme cheon2017homomorphic addressed machine learning use cases by supporting approximate arithmetic on encrypted data, prioritizing computational efficiency over exact precision. Brakerski, Gentry, and Vaikuntanathan’s BGV+ Scheme brakerski2014efficient introduced noise management optimizations to preserve model accuracy under encryption.
Xu et al. xu2019hybridalpha proposed HybridAlpha, a FL system combining differential privacy with secure aggregation to balance privacy and performance. While HybridAlpha reduces computational overhead compared to pure HE approaches, its reliance on generic privacy mechanisms leaves room for optimizations tailored to federated workflows.
2.2 Differential Privacy (DP) in Federated Learning
Differential Privacy (DP) reduces the risk of sensitive data exposure by perturbing model updates with controlled noise throughout the training phase. First formalized by Dwork et al. dwork2006calibrating , DP was later adapted to deep learning by Abadi et al. abadi2016deep , who developed the DP-SGD algorithm to protect training data. Mironov et al. mironov2017renyi refined these concepts using Rényi differential privacy, providing tighter privacy budget analysis for iterative machine learning processes. Despite these advancements, DP-based methods inherently degrade model utility due to noise injection, particularly in precision-sensitive applications like medical imaging.
2.3 Selective Encryption in Federated Learning
Selective encryption strategies aim to reduce computational overhead by encrypting only sensitive subsets of model parameters. Wu et al. wu2019towards demonstrated that selectively encrypting critical gradients preserves privacy while maintaining computational feasibility. Li et al. li2019privacy and Song et al. song2020privacy expanded this concept with adaptive parameter selection criteria, though their methods require careful tuning to avoid residual vulnerabilities.
Zhang et al. Zhang2020 introduced BatchCrypt, a HE framework that batches model updates to reduce communication and computation costs. While BatchCrypt improves efficiency over traditional HE methods, its encryption scope remains inflexible for dynamic FL scenarios. This paper extends prior work by integrating selective encryption with differential noise injection and bitwise scrambling, achieving enhanced privacy with reduced computational overhead.
2.4 Bitwise Scrambling for Federated Learning Security
@@
Bitwise scrambling enhances security by nonlinearly rearranging bits in partially encrypted data using cryptographic keys. Yang et al. yang2017bitwise applied scrambling to secure data transmissions against reconstruction attacks, demonstrating its effectiveness as a lightweight obfuscation layer. Halevi et al. halevi2019faster combined scrambling with HE to resist chosen-ciphertext attacks while maintaining computational efficiency.
The combination of selective encryption, differential privacy, and bitwise scrambling provides a multi-layered defense mechanism for FL systems. This integrated approach balances security robustness with practical efficiency, addressing the resource constraints inherent in distributed edge computing environments.
3 Motivation and Threat Model
3.1 Motivation
Applying FL to sensitive domains like healthcare and finance underscores a key issue: decentralized training by itself is insufficient to stop adversaries from uncovering sensitive patterns through exposed model updates. While FL avoids raw data centralization, empirical studies confirm that transmitted gradients retain sufficient information for model inversion attacks to reconstruct patient scans, financial transactions, or other identifiable records—even when training adheres to protocol.
Existing privacy mechanisms force practitioners into suboptimal trade-offs. FHE provides cryptographically rigorous confidentiality but imposes prohibitive computational latencies, rendering it impractical for real-time medical diagnostics or high-frequency trading systems. Differential privacy (DP), while computationally lightweight, irreversibly corrupts model updates with statistical noise, degrading diagnostic accuracy in oncology imaging or fraud detection below clinically/operationally viable thresholds.
Our FAS framework addresses this dichotomy through context-aware privacy stratification. By selectively encrypting only high-risk parameters (e.g., gradients correlating with identifiable features), applying tunable DP noise to less sensitive components, and augmenting both with bitwise scrambling, FAS maintains diagnostic-grade model utility while eliminating cryptographic overheads associated with full-model FHE. Crucially, this multi-tiered protection operates without requiring precomputed sensitivity masks or auxiliary pretraining phases, ensuring seamless integration into latency-constrained FL pipelines for MRI analysis, genomic prediction, and other precision-sensitive applications.
3.2 Threat Model
@@
FL systems are susceptible to privacy threats due to decentralized training and the exchange of model updates. A semi-honest attacker may adhere to the protocol while still trying to extract confidential information from observed or compromised data.
@@
Model updates shared between clients and the central server may be intercepted by attackers. Common attack vectors include gradient inversion to reconstruct training data, membership inference to verify specific data participation, and statistical analysis to infer private attributes. While SSL encryption secures transmissions, it does not prevent adversaries from analyzing transmitted updates.
@@
4 Proposed Model
This research presents FAS (Fast and Secure Federated Learning), a novel privacy-preserving scheme designed to reduce the computational cost of encryption while maintaining strong protection against privacy threats. The model integrates three lightweight techniques— selective encryption, differential noise addition, and bitwise scrambling along with HE—into the FL process. This multi-pronged approach enhances security without the heavy overhead typically associated with FHE or the accuracy degradation often introduced by differential privacy when used in isolation.
4.1 Selective Encryption
Instead of encrypting all model parameters, FAS applies HE only to a fixed percentage of the weights, selected uniformly. This subset is treated as the most sensitive part of the model. The encryption is performed directly on these weights using a homomorphic scheme, allowing the server to carry out aggregation without requiring decryption. This design choice retains the advantages of HE while reducing the time and resource demands substantially, achieving up to 90% reduction in overhead compared to full encryption.
4.2 Differential Noise
To reinforce privacy for the remaining (unencrypted) parameters, the model introduces controlled random noise. This noise follows a Laplace distribution and is calibrated using differential privacy principles. While FAS does not rely solely on differential privacy for security, this mechanism makes individual contributions harder to isolate, especially in low-sensitivity weights. By blending noise into less critical parameters, the model increases resistance to membership inference attacks with minimal impact on performance.
4.3 Bitwise Scrambling
FAS further obfuscates unencrypted data through bitwise scrambling. A lightweight cryptographic key is used to permute the bits of the non-encrypted parameters. This process disrupts patterns and statistical properties that could be exploited by adversaries. When combined with noise, scrambling ensures that unencrypted parts of the model are still resistant to reconstruction attacks—even in the presence of partial knowledge or auxiliary data.
4.4 Federated Learning Integration
FAS is designed to integrate directly into standard FL workflows. At each round, clients:
  1. 1.
    Train their local models on private data.
  2. 2.
    Encrypt a portion of the model weights.
  3. 3.
    Apply scrambling and noise to the rest.
  4. 4.
    Transmit the transformed updates to the server.
The server merges the incoming gradients using techniques like weighted averaging. Following this, a scrambling key is applied to the aggregated model before it is sent back to the clients, where it is decrypted and used for continued training.
4.5 Security and Efficiency Trade-Off
The proposed system achieves a balance between security and speed:
  • Security: FAS protects sensitive parameters through encryption while mitigating risks to the remaining ones through obfuscation. This layered defense shows comparable robustness to fully encrypted models in resisting attacks like model inversion and membership inference.
  • Efficiency: By reducing the encryption scope and avoiding pre-processing phases like sensitivity masking, FAS cuts down overhead significantly. It is particularly well-suited for settings with limited computational power.
By integrating selective encryption, noise injection, and scrambling, this approach offers a practical and deployable solution for secure FL in privacy-sensitive areas, including medical and financial domains.
4.6 Privacy-Preserving Techniques
4.6.1 Homomorphic Encryption
Homomorphic encryption (HE) is a cryptographic scheme that permits data processing to occur while the data remains encrypted, eliminating the need for decryption during computation.